A car and a bus set out a 2 p.m. from the same point, headed in the same direction. The average speed of the car is 30 mph slower than twice the speed of the bus. In two hours, the car is 20 miles ahead of the bus. Find the rate of the car.

Answer by checkley71(8403)

B=X SPEED OF THE BUS.
C=2X-30 SPEED OF THE CAR.
DISTANCE=R*T
SEEING AS THE DISTANCES OF THE BUS IS 20 MILES LESS WE HAVE:
2(2X-30)=2X+20 NOTE WE HAVE TO ADD 20 MILES TO THE BUS DISTANCE TO MAKE THEM EQUAL.
4X-60=2X+20
4X-2X=20+60
2X=80
X=80/2
X=40 MPH FOR THE BUS.
C=2*40-30
C=80-30
X=50 PMH FOR THE CAR.
PROOF:
2*50=2*40+20
100=80+20
100=100